# Uncertainties from metrology in the integrated luminosity measurement with the updated design of a detector at CEPC

### v3

@@ 1 Introduction
In this paper, we review the potential impact of locating the luminometer around the z-axis on uncertainties associated with measuring integrated luminosity. These uncertainties stem from mechanical factors like positioning and alignment of the luminometer and the fact that the beam properties are known within some margins. These effects are collectively referred to as metrology. Additionally, we examined the performance of different Bhabha counting methods, including the LEP-style asymmetric counting OPAL , typically utilized to mitigate systematic biases arising from asymmetries between the left and right detector halves.
The paper is organized as follows: in Section 2 we discuss the current layout of the very forward region of CEPC. In Section 3 we review the impact of detector manufacturing, positioning and alignment (Section 3.1), as well as the impact of uncertainties of the knowledge of the beam properties (asymmetry of beam energies, beam energy spread - BES and beam synchronization) in Section 3.2. The concluding discussion is given in Section 4.
@@ 2 Forward region at CEPC
The choice of technology of the luminometer at CEPC has not been finally determined yet. However, since this study is based on simulation of counting of low-angle Bhabha scattering without any other assumption but the angular acceptance of the luminometer, it is of no relevance for this study. If one would perform a more realistic simulation including detector response to signal and background, at least an energy cut on electron and positron tracks should be required to separate Bhabha events from the background. In such a case, effects like bias and resolution of energy measurement as well as the uncertainties from calibration of the luminometer should be included as additional sources of systematic uncertainty. The same holds if one includes other selection criteria based on polar and azimuthal angles of a Bhabha candidate.
@@ 3 Integrated luminosity measurement and systematic uncertainties
where $\alpha_{em}$ is the QED constant and $\theta$ the polar angle of emitted Bhabha particles. Uncertainty of $\mathrm{10^{-4}}$ of the cross-section implies that the available center-of-mass energy should be known with the absolute uncertainty of $\Delta(\sqrt{s})\leq$ 5 MeV. There is an ongoing work at future Higgs factories nesto to check the feasibility of $\sqrt{s}$ determination using di-muon production reconstructed in the central tracker. However, the required precision seems to be beyond the experimental reach, but, as discussed in Stahl , many s-channel processes far from the resonance will have the same dependence of the cross-section on the center-of-mass energy, leading to effective cancellation of the uncertainties in the cross-section measurements. It is up to each individual cross-section measurement analysis to determine if uncertainty of the integrated luminosity measurement caused by the uncertainty of $\sqrt{s}$ is an issue.
@@
Further, feasibility and requirements for $\mathrm{10^{-4}}$ precision of counting at the $Z^{0}$ pole will be discussed, considering that the uncertainty of each individual systematic effect contributes to the integrated luminosity at a level of $\mathrm{10^{-4}}$. In JINST , at one side of the luminometer the Bhabhas are counted in the full fiducial volume of the luminometer, while at the other side the inner radial acceptance of the fiducial volume is shrunken by 1 mm. In this study, the outer radial acceptance is also shrunken by 1 mm, as it has been done at OPAL OPAL , randomly to the left (L) and right (R) sides of the luminometer, on event by event basis. Described way of counting is called asymmetrical throughout the paper. For symmetric event selection the Bhabhas are counted in the full fiducial volumes of the both sides of luminometer. As will be discussed further, different detector positioning (s-axis and z-axis) interfere with a way of counting in a nontrivial way.
@@ 3.1 Uncertainties from mechanics and positioning
Considered detector-related uncertainties originating from manufacturing, positioning and alignment of the luminometer are, as in JINST :
@@
    Maximal absolute uncertainty of the distance between left and right sides of the luminometer along the z-axis, $\Delta l$, where both halves are shifted equidistantly for ±$\pm\frac{\Delta l}{2}$ with respect to the interaction point,
@@
    RMS of the Gaussian distribution of luminometer fluctuations with respect to the IP, caused by vibrations and thermal stress, in radial, $\sigma_{x_{IP}}$, and axial direction, $\sigma_{z_{IP}}$.
@@
Figures 4-8 illustrate effects 1-4 respectively, for different ways of counting and different positions of the luminometer placed on the s- and z-axis. Values in histograms shown on Figures 4(a)-8(a) are based on the data from Figures 4(b)-8(b), obtained from the local fit in the region of interest. If the distribution on (b) was not symmetrical for positive and negative variations of the parameter, the more restrictive limit was considered.
@@
Figures 4-8 show that for the detector placed at the z-axis asymmetric counting performs similarly as counting in the full fiducial volume, whereas with the detector placed at the s-axis the difference between the two ways of counting is quite significant. On Figure 4, which illustrates the impact of $\Delta r_{in}$, it can be seen that the counting uncertainties with the detector placed at the z-axis are similar to uncertainties when counting in the full fiducial volume with the detector placed at the s-axis. For $\sigma_{r}$ and $\sigma_{z_{IP}}$, the requirements are slightly more relaxed for the luminometer placed at the z-axis than with the asymmetric counting with the luminometer placed at the s-axis, as shown in Figures 5 and 8, whether for $\sigma_{x_{IP}}$ it’s the other way around. For $\Delta l$, the uncertainties are the same for both positions of the luminometer and for both ways of counting, as illustrated in Figure 6. However, one has to keep in mind that the results shown in this paper are based on distributions of mean values, each obtained from the simulated signal sample of $2\cdot 10^{7}$ events, with intrinsic relative statistical uncertainty of $\sim 2\cdot 10^{-4}$. Therefore, the extracted precision limits rather reflect the order of magnitude required to control certain mechanical or beam properties at the level of $10^{-4}$ than the exact values of detector or beam parameters.
@@
Figure 4: (a) Variations of the inner aperture (counting volume), $\Delta r_{in}$, for two different ways of counting: in the full fiducial volume (S) and the LEP-style (AS), for detector placed on the s- and z-axis; (b) Count uncertainty as a function of $\Delta r_{in}$ for two different ways of counting (S and AS), with the detector placed at the s- and z-axis.
Figure 5: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for the RMS of the Gaussian spread of measured radial shower position with respect to the true one, $\sigma_{r}$. (b) Count uncertainty as a function of $\sigma_{r}$ for two different ways of counting (S and AS), with the detector placed at the s- and z-axis.
@@
Figure 6: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for maximal absolute uncertainty, $\Delta l$, of the longitudinal distance between left and right luminometer sides; (b) Count uncertainty as a function of $\Delta l$ for two different ways of counting (S and AS), with the detector placed at the s- and z-axis.
@@
Figure 7: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis against the RMS of a Gaussian distribution of radial fluctuations of the luminometer position with respect to the IP, $\sigma_{x_{IP}}$; (b) Count uncertainty as a function of $\sigma_{x_{IP}}$ for two different ways of counting (S and AS), with the detector placed at the s- and z-axis.
@@
Figure 8: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis against the RMS of the Gaussian distribution of axial fluctuations of the luminometer position with respect to the IP, $\sigma_{z_{IP}}$; (b) Count uncertainty as a function of $\sigma_{z_{IP}}$ for two different ways of counting (S and AS), with the detector placed at the s- and z-axis.
@@ 3.2 Beam related uncertainties
Several systematic uncertainties may arise from uncertainties of the properties of the beams and their delivery to the interaction point. The list of CEPC beam parameters addressed in this study is given in Table 1 Gao .
@@
Here we consider, again similarly to JINST :
@@
    maximal time shift in beam synchronization ($\Delta\tau_{SY}$), determined from $\Delta z_{IP}^{SY}$, which may cause the IP longitudinal displacement $\Delta z_{IP}^{SY}$.
In the same manner as in the previous chapter, results for effects 1-4 are illustrated on Figures 9-12 and the upper limits, together with the upper limits of mechanical parameters from Section 3.1, are summarized in Table 2. Each effect is considered to individually contribute to the the relative uncertainty of the integrated luminosity as $1\cdot 10^{-4}$.
Figure 9: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for maximal bias of energy of one beam with respect to the other, $\Delta E$; (b) Loss of the Bhabha count in the luminometer due to the longitudinal boost of the center-of-mass frame, $\beta_{z}(\Delta E)$, for two different ways of counting (S and AS), with the detector placed at the s- and z-axis.
Figure 10: (b) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for maximal RMS of the BES Gaussian distribution, corresponding to the relative systematic uncertainty of $\mathcal{L}$ of $10^{-4}$, $\sigma_{E_{BS}}$; (b) Loss of the Bhabha count in the luminometer due to longitudinal boost of the Bhabha center-of-mass frame caused by BES, $\beta_{z}(\sigma_{E_{BS}})$, for two different ways of counting (S and AS), with the detector placed at the s- and z-axis.
Figure 11: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for maximal radial IP position displacement with respect to the luminometer ($\Delta x_{IP}^{BS}$); (b) Count uncertainty as a function of $\Delta x_{IP}^{BS}$ for two different ways of counting (S and AS), with the detector placed at the s- and z-axis.
Figure 12: (a) A comparison of results between the luminometer placed at the s-axis and luminometer placed at the z-axis for maximal axial IP position displacements with respect to the luminometer ($\Delta z_{IP}^{BS}$); (b) Count uncertainty as a function of $\Delta z_{IP}^{BS}$ for two different ways of counting (S and AS), with the detector placed at the s- and z-axis.
@@
$\Delta\tau$ (ps) 0.08 3 3 3
As can be seen from Table 2, mechanical precision of the luminometer’s inner aperture seems to be the major challenge, since the Bhabha cross-section is scaling with the polar angle as $\sigma\sim 1/\theta^{3}$. The longitudinal boost of the Bhabha center-of-mass frame with respect to the laboratory frame, $\beta_{z}$, may be caused by any difference in beam energies, $\delta E$ ($\beta_{z}=2\cdot\delta E/\sqrt{s}$), in the form of bias of one beam with respect to the other, $\Delta E$, or as a random asymmetry caused by the beam energy spread or by other radiative losses. All limits of parameters given in Table 2 are derived from distributions of mean values in Figures 4(b) - 12(b), without considering statistical uncertainties from the limited size of the simulated samples. From Table 2 it can be seen that in terms of the beam-related parameters precision, displacement of the luminometer from the s-axis to the z-axis will not impact the luminosity precision significantly, while some dependencies between observables, especially $\Delta x_{IP}^{BS}$, $\Delta z_{IP}^{SY}$ (Figures 11 and 12) will be modified due to loss of spacial symmetries of Bhabha events.
@@ 4 Conclusion
Control of the luminometer inner radius at the micrometer level at the $Z^{0}$ pole seeems to be the most demanding requirement for the detector manufacturing, regardless of the luminometer’s positioning (on the s-axis or on the z-axis). Any eventual change in the luminometer design towards smaller polar angle coverage will require control of the luminometer inner radius even below micrometer precision, according to the dependence of the Bhabha cross-section on the polar angles of the scattered particles. However, it is important to note that these results are obtained under assumption that the change of the luminometer’s inner radius corresponds to exactly the same change of the inner radius of the luminometer’s fiducial volume, which needs to be approved. Also, control of the asymmetrical bias in beam energies might be challenging, while the beam energy spread can be relaxed significantly beyond the current RMS of 18 MeV at the $Z^{0}$ pole.
From the point of view of experimental input for the Bhabha cross-section calculations, center-of-mass energy should be known at the $\mathrm{10^{-4}}$ level, which requires further studies in terms of feasibility of such a precision. This might be an issue for the cross-section measurements on the slope of a production threshold.
Although the simulated Bhabha samples are relatively small to be exact on boundaries of the considered systematics, distribution of central values presented in this study confirms that no effect seems to be more critical for precision measurement of the integrated luminosity if the luminometer is placed around the z-axis than if it is placed around the outgoing beams. The LEP-style asymmetric counting, however, will not be effective for the luminometer placed on the z-axis to cope with the effects arising from the left-right asymmetry.

### v4

@@ 1 Introduction
In this paper, we review the potential impact of positioning the luminometer around the z-axis on systematic uncertainties associated with measuring integrated luminosity. These uncertainties stem from mechanical factors like positioning and alignment of the luminometer, as well as from the fact that beam properties are known within some margins. These effects are collectively referred to as metrology. Additionally, we examined the performance of different Bhabha counting methods, including the LEP-style asymmetric counting LEP1 , typically utilized to mitigate systematic biases arising from asymmetries between the left and right detector halves.
The paper is organized as follows: in Section 2 we discuss the current layout of the very forward region of CEPC. In Section 3 we review the impact of detector manufacturing, positioning and alignment (Section 3.1), as well as the impact of uncertainties of the beam properties (asymmetry of beam energies, beam energy spread - BES, IP position and beam synchronization) in Section 3.2. The concluding discussion is given in Section 4.
@@ 2 Forward region at CEPC
The choice of technology of the luminometer at CEPC has not been finally determined yet. However, since this study is based on simulation of counting of low-angle Bhabha scattering without any other assumption but the angular acceptance of the luminometer, it is of no relevance for this study. If one would perform a more realistic simulation including detector response to signal and background, at least an energy cut on electron and positron tracks should be required to separate Bhabha events from the background. In such a case, effects like bias and resolution of energy measurement as well as the uncertainties from calibration of the luminometer should be included as additional sources of systematic uncertainty. The same holds if one includes other selection criteria based on polar and azimuthal angle of Bhabha candidates.
@@ 3 Integrated luminosity measurement and systematic uncertainties
where $\alpha_{em}$ is the QED constant and $\theta$ the polar angle of emitted Bhabha particles. Uncertainty of $\mathrm{10^{-4}}$ of the cross-section implies that the available center-of-mass energy should be known with the absolute uncertainty of $\Delta(\sqrt{s})\leq$ 5 MeV. There is an ongoing study nesto at future Higgs factories to check the feasibility of $\sqrt{s}$ determination using di-muon production reconstructed in the central tracker. However, the required precision seems to be beyond the experimental reach, but, as discussed in Stahl , many s-channel processes far from the resonance will have the same dependence of the cross-section on the center-of-mass energy, leading to effective cancellation of the uncertainties in the cross-section measurements. It is up to each individual cross-section measurement analysis to determine if uncertainty of the integrated luminosity measurement caused by the uncertainty of $\sqrt{s}$ is an issue.
@@
Further, feasibility and requirements for $\mathrm{10^{-4}}$ precision of counting at the $Z^{0}$ pole will be discussed, considering that the uncertainty of each individual systematic effect contributes to the integrated luminosity at a level of $\mathrm{10^{-4}}$. In JINST , at one side of the luminometer the Bhabhas are counted in the full fiducial volume of the luminometer, while at the other side the inner radial acceptance of the fiducial volume is shrunken by 1 mm. In this study, the outer radial acceptance is also shrunken by 1 mm. This is performed randomly to the left (L) and right (R) sides of the luminometer, on event by event basis, as it has been done at ALEPH LEP1 . Described way of counting is called asymmetrical (AS) throughout the paper. For symmetric (S) event selection the Bhabhas are counted in the full fiducial volumes of the both sides of luminometer. As will be discussed further, different detector positioning (s-axis and z-axis) interfere with a way of counting in a nontrivial way.
@@ 3.1 Uncertainties from mechanics and positioning
Considered detector-related uncertainties originating from manufacturing, positioning and alignment of the luminometer are:
@@
    RMS of the Gaussian distribution of luminometer fluctuations with respect to the IP, caused by vibrations and thermal stress, in radial, $\sigma_{x_{IP}}$, and axial direction, $\sigma_{z_{IP}}$,
@@
    Maximal absolute uncertainty of the distance between left and right sides of the luminometer along the z-axis, $\Delta l$, where both halves are shifted equidistantly for ±$\pm\frac{\Delta l}{2}$ with respect to the interaction point.
@@
Figures 4-7 illustrate effects 1-4 respectively, for different ways of counting and different positions of the luminometer placed on the s- and z-axis. Values in histograms shown on Figures 4(b)-7(b) are based on the data from Figures 4(a)-7(a), obtained from the local fit in the region of interest. If the distribution on (a) was not symmetrical for positive and negative variations of the parameter, the more restrictive limit was considered.
@@
Figures 4-7 show that for the detector placed at the z-axis asymmetric counting performs similarly as counting in the full fiducial volume, whereas with the detector placed at the s-axis the difference between the two ways of counting is quite significant. In Figure 4, which illustrates the impact of $\Delta r_{in}$, it can be seen that the requirement on the precision of inner radius of the counting volume is relaxed with the detector placed at the s-axis. This is due to the fact that asymmetric counting allows $\Delta r_{in}$ to modify one side of the luminometer, while the other side of the counting volume is shrunken, compensating for the uncertainty $\Delta r_{in}$. This compensation works only if signal hits are L-R symmetrical, which is the case for the detector placed at the s-axis, and effect itself introduces L-R asymmetry. When the detector is placed at the z-axis, the symmetry of signal hits is lost (Figure 2) and, therefore, in this case asymmetric counting is no better than the symmetric counting in the full FV. As illustrated for $\sigma_{r}$ (Figure 5), $\sigma_{x_{IP}}$ (Figure 6) and $\sigma_{z_{IP}}$ (Figure 7), asymmetric counting works significantly better than symmetric for the detector placed at the s-axis and doesn’t make a difference for the detector placed at the z-axis. If an effect does not introduce L-R asymmetry, like $\Delta l$ illustrated in Figure 8, there would be no difference if counting was performed in the full FV on both sides of the luminometer, or in the asymmetric way. One has to keep in mind that the results shown in this paper are based on distributions of mean values, each obtained from the simulated signal sample of $2\cdot 10^{7}$ events, with intrinsic relative statistical uncertainty of $\sim 2\cdot 10^{-4}$. Therefore, the extracted precision limits rather reflect the order of magnitude required to control certain mechanical or beam properties than to be the exact values of detector or beam parameter uncertainties.
@@
Figure 4: (a) Count uncertainty as a function of $\Delta r_{in}$ for two different ways of counting: in the full fiducial volume (S) and the LEP-style (AS), with the detector placed at the s- and z-axis; (b) Variations of the inner aperture of the counting volume, $\Delta r_{in}$, for two different ways of counting (S and AS), for detector placed on the s- and z-axis. Precision limits are taken from the distributions given in (a) for $\Delta N/N$ uncertainty of $10^{-4}$.
@@
Figure 5: (a) Count uncertainty as a function of $\sigma_{r}$ for two different ways of counting (S and AS), with a detector placed at the s- and z-axis. (b) A comparison of results for $\sigma_{r}$ for luminometer placed at the s-axis and at the z-axis. Precision limits are taken from the distributions given in (a) for $\Delta N/N$ uncertainty of $10^{-4}$.
@@
Figure 6: (a) Count uncertainty as a function of $\sigma_{x_{IP}}$ for two different ways of counting (S and AS), with a detector placed at the s- and z-axis; (b) A comparison of results for $\sigma_{x_{IP}}$ for luminometer placed at the s-axis and at the z-axis. Precision limits are taken from the distributions given in (a) for $\Delta N/N$ uncertainty of $10^{-4}$.
@@
Figure 7: (a) Count uncertainty as a function of $\sigma_{z_{IP}}$ for two different ways of counting (S and AS), with a detector placed at the s- and z-axis; (b) A comparison of results for $\sigma_{z_{IP}}$ for luminometer placed at the s-axis and at the z-axis. Precision limits are taken from the distributions given in (a) for $\Delta N/N$ uncertainty of $10^{-4}$.
Figure 8: (a) Count uncertainty as a function of $\Delta l$ for two different ways of counting (S and AS), with a detector placed at the s- and z-axis; (b) A comparison of results for $\Delta l$ for luminometer placed at the s-axis and at the z-axis. Precision limits are taken from the distributions given in (a) for $\Delta N/N$ uncertainty of $10^{-4}$.
Figure 9: (a) Loss of the Bhabha count in the luminometer due to the longitudinal boost of the center-of-mass frame, $\beta_{z}(\Delta E)$, for two different ways of counting (S and AS), with a detector placed at the s- and z-axis; (b) A comparison of results for $\Delta E$ for luminometer placed at the s-axis and at the z-axis. Precision limits are taken from the distributions given in (a) for $\Delta N/N$ uncertainty of $10^{-4}$.
Figure 10: (b) Loss of the Bhabha count in the luminometer due to longitudinal boost of the Bhabha center-of-mass frame caused by BES, $\beta_{z}(\sigma_{E_{BS}})$, for two different ways of counting (S and AS), with a detector placed at the s- and z-axis; (b) A comparison of results for $\sigma_{E_{BS}}$ for luminometer placed at the s-axis and at the z-axis. Precision limits are taken from the distributions given in (a) for $\Delta N/N$ uncertainty of $10^{-4}$.
Figure 11: (a) Count uncertainty as a function of $\Delta x_{IP}^{BS}$ for two different ways of counting (S and AS), with a detector placed at the s- and z-axis; (b) A comparison of results for $\Delta x_{IP}^{BS}$ for luminometer placed at the s-axis and at the z-axis. Precision limits are taken from the distributions given in (a) for $\Delta N/N$ uncertainty of $10^{-4}$.
Figure 12: (a) Count uncertainty as a function of $\Delta z_{IP}^{BS}$ for two different ways of counting (S and AS), with a detector placed at the s- and z-axis; (b) A comparison of results for $\Delta z_{IP}^{BS}$ for luminometer placed at the s-axis and at the z-axis. Precision limits are taken from the distributions given in (a) for $\Delta N/N$ uncertainty of $10^{-4}$.
@@ 3.2 Beam related uncertainties
Several systematic uncertainties may arise from uncertainties of the properties of the beams and their delivery to the interaction point. The list of main CEPC beam parameters is given in Table 1 Gao .
@@
Here we consider:
@@
    maximal time shift in beam synchronization ($\Delta\tau_{SY}$), causing the IP longitudinal displacement $\Delta z_{IP}^{SY}$.
In the same manner as in the previous chapter, results for effects 1-4 are illustrated in Figures 9-12 and the upper limits, together with the upper limits of mechanical parameters from Section 3.1, are summarized in Table 2. All limits of parameters given in Table 2 are derived from distributions of mean values in Figures 4(a) - 12(a), without considering statistical uncertainties from the limited size of the simulated samples. Each effect is considered to individually contribute to the the relative uncertainty of the integrated luminosity as $1\cdot 10^{-4}$.
@@
$\Delta\tau$ (ps) 0.2 7 7 7
As can be seen from Table 2, mechanical precision of the inner aperture of the counting volume seems to be the major challenge, since the Bhabha cross-section is scaling with the polar angle as $\sigma\sim 1/\theta^{3}$. However, asymmetric counting applicable at the s-axis relaxes this requirement, while for the z-axis the inner radius of the counting volume should be controlled with precision $\leqslant 1$ $\mathrm{\mu m}$. The longitudinal boost $\delta E$ of the Bhabha center-of-mass frame with respect to the laboratory frame, $\beta_{z}$, may be caused by any difference in beam energies ($\beta_{z}=2\cdot\delta E/\sqrt{s}$), either as a bias of energy of one beam with respect to the other, $\Delta E$, or as a random asymmetry in beam energies caused by the beam energy spread or by other radiative losses. While the beam energy spread at CEPC is not an issue for integrated luminosity measurement at the $Z^{0}$ pole, bias in beam energies, $\Delta E$, can be tolerated up to $\mathrm{7\>MeV}$ ($1.5\cdot 10^{-4}$ of the nominal beam energy). From Table 2 it can be seen that displacement of the luminometer from the s-axis to the z-axis will not impact the luminosity precision significantly, while some dependencies will be modified due to loss of spacial symmetries of Bhabha events.
@@ 4 Conclusion
Control of the luminometer inner radius at the micrometer level at the $Z^{0}$ pole seeems to be the most demanding requirement for the detector manufacturing. Any additional change in the luminometer design towards smaller polar angle coverage will require control of the luminometer inner radius even below micrometer precision, according to the dependence of the Bhabha cross-section on the polar angles of the scattered particles. However, it is important to note that these results are obtained under assumption that the change of the luminometer’s inner radius corresponds to exactly the same change of the inner radius of the luminometer’s fiducial volume, which needs to be approved. Also, control of the asymmetrical bias in beam energies might be challenging ($1.5\cdot 10^{-4}$), while the beam energy spread can be relaxed significantly beyond the current RMS of 18 MeV at the $Z^{0}$ pole. Existing techniques like the Frequency Scanning Interferometry (FSI) allow positioning control of the luminometer at the micrometer level Daniluk , thus the mechanical effects from metrology should not be issue for the integrated luminosity measurement.
From the point of view of experimental input for the Bhabha cross-section calculations, center-of-mass energy should be known at the $\sim{10^{-4}}$ level, which requires further studies in terms of feasibility of such a precision. This might be an issue for the cross-section measurements on the slope of a production threshold.
Although the simulated Bhabha samples are relatively small to be exact on boundaries of the considered systematics, distribution of central values presented in this study confirms that no effect seems to be more critical for precision measurement of the integrated luminosity if the luminometer is placed around the z-axis than if it is placed around the outgoing beams, except for the inner aperture of the counting volume that can be relaxed with the asymmetric counting at the s-axis. The LEP-style asymmetric counting, however, will not be effective for the luminometer placed on the z-axis to cope with the effects arising from the left-right asymmetry.